Focus Motion Failures

Saturday, October 19, 2019 9:04 AM

Problem(s) Encountered:

Had an issue with the telescope focus tonight. I noticed it wasn't maintaining the focus I set at the beginning of the night (was supposed to be 5275, I caught it at 5299). Couldn’t move it my giving it a focus value in the  Misc  window. I power cycled the TCS computer and the chassis, but still it wouldn't budge through the  Misc  window. Throughout these attempts to adjust the focus in this way I could hear the motors running, but to no avail. I finally tried to move focus with the hand paddle at its fastest speed which worked. I was able to move the focus with the  Misc  window after that. Lost about 20 minutes.


This problem occurred again on a later run.  Not sure what is causing these longstanding failures with focus response.  Have a query in with DFM and will see where it goes.That said, it seems that just waiting seems to resolve the issue (I know this is not a satisfactory solution).  Everything else appears to be red herrings.